# 7-3-1 Hongo, Bunkyo-ku, Tokyo 113-0033, Japan

### v3

@@ Abstract
We apply the background charged bosonic free-field approach to the rational principal quantum Drinfeld-Sokolov (QDS) $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ minimal models with boundaries, where $g$ is a simple bosonic Lie algebra. We express their Ishibashi states using the free-field Ishibashi states, and calculate disk two-point correlation functions using the free-field vertex operators with intertwiner insertions. The integral results are obtained by repeatedly applying the Euler-type integral expression and the Taylor expansions of Lauricella’s hypergeometric functions $F_{D}^{(n)}$. We also discuss the background free-field approach to the QDS $\mathcal{W}$ algebras related to a semi-simple Lie algebra $g$, and apply a coset free-field resolution to express diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states.
@@ 1 Introduction
Two-dimensional conformal field theories (CFT2) have infinite-dimensional local conformal symmetries, described by a local conformal algebra $\mathcal{A}\otimes\widebar{\mathcal{A}}$. $\mathcal{A}$ and $\widebar{\mathcal{A}}$ are infinite-dimensional Lie algebras, called the chiral and anti-chiral algebras, respectively Belavin:1984vu ; Moore:1988qv . The infinite-dimensional local conformal symmetries allow us to solve the correlation functions in some theories exactly, and free-field approaches are efficient methods. We claim that a CFT2 admits a free-field approach if the following three conditions are satisfied Bouwknegt:1990wa .
@@
There is a wide range of RCFTs that are known to admit free-field approaches. Two canonical types of theories with free-field approaches are
We apply the background-charged free-field approach to rational QDS $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ ($g$ being simple Lie algebras), including free-field expressions of Ishibashi states, and detailed example calculations of disk correlation functions using the free-field vertex operators, with insertions of $\mathcal{W}$ intertwiners. The method to calculate the correlation functions is a direct generalization of the celebrated Coulomb gas formalism of the Virasoro minimal models Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi . We show that the integral expressions of this type are solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type of differential equations of the QDS $\mathcal{W}$ minimal models Belavin:1984vu . The crucial techniques in performing the integrals obtained from the free-field approaches are:
@@
Divergent terms can arise in the process. However, we have shown that the integrals expressions are solutions to the BPZ type of differential equations, it is necessary to include them. We give sample calculations of disk two-point functions in some diagonal and charge-conjugated Virasoro minimal models, $\mathcal{W}_{3}$ minimal models, the charge-conjugated $\mathcal{W}_{4}(7,4)$ model, and the $\mathcal{W}\big{[}\widehat{G}_{2}\big{]}(6,7)$ model. The calculations of the correlation functions with higher points and in other rational $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ models can be obtained from analogous procedures in principle.
In the appendices, we discuss the background-charged free-field approach to the QDS $\mathcal{W}$ algebras related to semi-simple Lie algebras. The free-field approach reveals that each simple part of a semi-simple QDS $\mathcal{W}$ algebra decouples. The chiral characters, modular properties, fusion rules, and correlation functions of the semi-simple $\mathcal{W}$ theory are a multiplication of those of the simple parts. Finally, we discuss applying a coset resolution Bouwknegt:1990fb ; Bouwknegt:1990wa to express the diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states Goddard:1984vk ; Goddard:1986ee ; Mathieu:1990dy ; Bouwknegt:1992wg ; Arakawa:2018iyk .
@@ 2.1 Preliminaries
For each element $w\in W(g)$, we can define its length $l(w)$ as the minimum possible number of elements in its simple Weyl reflection decompositions. There is a unique longest element $w_{l}\in W(g)$ and the charge conjugation of $\lambda$ is defined as $U\lambda:=-w_{l}\lambda$. For the $B_{r}$, $C_{r}$, $D_{r}$ even series, $E_{7,8}$, $F_{4}$, and $G_{2}$ the charge conjugation $U=-w_{l}=I$. For the $A_{r}$ series, $U$ reverse the order of all Dynkin labels. For the $D_{r}$ odd series, $U$ reverse the final two Dynkin labels $\lambda_{r-1}\leftrightarrow\lambda_{r}$. For $E_{6}$, $U$ reverse two pairs of Dynkin labels $\lambda_{1}\leftrightarrow\lambda_{6}$ and $\lambda_{3}\leftrightarrow\lambda_{5}$. We define two types of shifted Weyl actions as
@@
Dominant and codominant weight vectors are defined analogously to the finite cases, with the addition of the extra zeroth dimension. The set of level-$k$ dominant and codominant $\widehat{g}(k)$ weight vectors are denoted by $\widehat{P}_{+}^{k}$ and $\widehat{P}_{+}^{\vee k}$ respectively. The principal admissible weight vectors are defined in Frenkel:1992ju . We are only interested in the admissible $\widehat{g}(k)$ weights taking the form of
@@
Define the screening operators of a chiral algebra $\mathcal{A}$ as the free-field vertex operators whose OPEs with free-field chiral generators are either a regular or a total derivative singular term. From the definition, we conclude that a closed contour integral over an $\mathcal{A}$ screening operator is an $\mathcal{A}$ intertwiner.
The free-field approach to the simple KM $\widehat{g}(k)$ algebras we used here is the Wakimoto-type, which is realized from $r$ background-charged bosonic fields $\phi^{i}$ and $\big{|}\Delta_{+}\big{|}$ bosonic $\beta\gamma$ ghost systems Wakimoto:1986gf ; Bouwknegt:1990aa ; Bouwknegt:1990wa . Our focus is to present a free-field resolution conjecture for dominant and admissible modules $\widetilde{L}_{\widehat{\Lambda}}$ using the $\phi^{i}\beta^{\alpha}\gamma^{\alpha}$ Fock space modules, since the Fock space resolution of irreducible QDS $\mathcal{W}$ minimal modules are induced from them. Consider a simple KM algebra at positive integer level $k\in\mathbb{Z}_{+}$, where the chiral modules are the irreducible dominant modules $\widetilde{L}_{\widehat{\Lambda}}$, $\widehat{\Lambda}\in\widehat{P}^{k}_{+}$. The screening operators of this theory take the form of Bouwknegt:1990wa
@@
Let $C_{(\Lambda^{(+)},\Lambda^{(-)})}=(F_{(\Lambda^{(+)},\Lambda^{(-)})}^{\phi^{i}},d)$ be the BRST image of the $\widehat{g}(k)$ admissible resolution complex $\widetilde{C}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$, where $d$ are $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ intertwiners. The BRST image of an admissible (and dominant) $\widehat{g}(k)$ modules is either zero or irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ of the $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ algebras Frenkel:1992ju . Hence, the zeroth cohomology of the complex $(F_{(\Lambda^{(+)},\Lambda^{(-)})},d)$ is either trivial or being isomorphic to the irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ Bouwknegt:1991gf .
@@
$\Big{\{}L_{(\Lambda^{(+)},\Lambda^{(-)})}\quad|\quad\widehat{\Lambda}^{(+)}\in\widehat{P}^{p-h^{\vee}}_{+},\>\widehat{\Lambda}^{(-)}\in\widehat{P}^{p^{\prime}-h}_{+}\Big{\}},$ (20)
@@
$(L_{n}-\widebar{L}_{-n})|b\rangle_{\Omega}=\big{[}W_{n}^{i}-\Omega(-1)^{h_{i}}\widebar{W}^{i}_{-n}\big{]}|b\rangle_{\Omega}=0.$ (25)
For each $\mathcal{A}$ representation $L_{i}$ (not necessarily irreducible), the solutions to the conditions (25) are the $\Omega$-twisted Ishibashi states $|i,b\rangle\rangle_{\Omega}$ Ishibashi:1988kg ; Onogi:1988qk
$|i\rangle\rangle_{\Omega}=\sum_{N=0}^{\infty}|i,N\rangle\otimes V_{\Omega}U|i,N\rangle,$ (26)
where $|i,N\rangle$ denotes the orthonormal basis of $L_{i}$. $U$ is an anti-unitary isomorphism mapping a representation to its charge conjugation. $V_{\Omega}$ is a unitary isomorphism induced from $\Omega$ such that $V_{\Omega}:\>L_{i}\to L_{\omega^{-1}(i)}$ Recknagel:1997sb ; Recknagel:2013uja . The physical boundary states need to be constrained by the following condition: the open-sector Hilbert space between two boundaries must take the form of
@@
One simple type of solution satisfying this constraint is the Cardy boundary states, where the boundary states have the same labeling as the modules in the Kac-table, with the celebrated $n=N$ relation is imposed Cardy:1989ir . The forms of Cardy boundary states are
$|b_{i}\rangle_{\Omega}=\sum_{j}\frac{S_{ij}}{\sqrt{S_{0j}}}|j\rangle\rangle_{\Omega}.$ (28)
@@ 2.2 Universal results
$|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{\Omega}=\sum_{\widehat{w}\in\widehat{W}}\theta_{\widehat{w}}|F_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\rangle\rangle_{\Omega^{\prime}},$ (29)
@@
(33) means that all the $W^{i}$ gluing conditions are satisfied if the Virasoro gluing condition, and the anti-holomorphic momentum satisfy the required conditions.
@@
We conjecture that $w^{i}$ eigenvalues are the standard symmetric polynomial of order $i$ of the variables $\theta_{j}=(\Lambda+\alpha_{+}\rho+\alpha_{-}\rho^{\vee},\epsilon_{j})$, where $\{\epsilon_{j}\}$ is the set of weight of vector representation of $g$. This conjecture is a direct generalization of the $w^{3}$ eigenvalue shown in Bouwknegt:1992wg . The $w^{i}$ eigenvalues changed under the action $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ as $w^{i}\to(-1)^{h^{i}}w^{i}$. We show that this is how $w^{i}$ eigenvalues change under charge conjugation. The action of charge conjugation $U$ on the KM $\widehat{g}(k)$ currents as the conjugation automorphism $H^{i}(z)\to-H^{i}(z)$, $E^{\pm\alpha}(z)\to E^{\mp\alpha}(z)$, combined with the fact that $W^{i}(z)$ generators are obtained from the Hamiltonian reductions of the KM currents with $h^{i}$ order. Hence, the $w^{i}$ eigenvalues tramsform under charge conjugation as
$U:\quad w^{i}\to(-1)^{h^{i}}w^{i}.$ (35)
Hence, we can understand that the shift $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ is a form of charge conjugation in the free-field realization.
@@
$(W^{i}_{0}-(-1)^{h^{i}}\widebar{W}^{i}_{0})|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{I}=0,$ (36)
@@
The method we use to calculate the $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ correlation functions, is a direct generalization of the Coulomb gas formalism Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi , where the sphere correlation functions are calculated from free-field vertex operators with insertions of intertwiners to satisfy the neutrality condition. We prove the validity of this method by showing that the integral expressions obtained are the solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type differential equations when the neutrality conditions are satisfied Belavin:1984vu .
@@
$\langle\chi(z)\cdots\rangle=D\langle\mathcal{O}(z)\cdots\rangle=0.$ (38)
The chiral algebra intertwiners map $\mathcal{A}$ primary vectors to $\mathcal{A}$ primary vectors, hence, we could write any singular vector as $|\chi\rangle=Q^{\chi}|v\rangle$, where $Q^{\chi}$ is an intertwiner and $|v\rangle$ is a primary vector Thorn:1984sn ; Bouwknegt:1992wg . This allow us to replace the singular operator by a commutator $Q^{\chi},\mathcal{O}(z)]$, for some local operator $\mathcal{O}(z)$. Write the intertwiner as contour integral over a screening operator, we obtain
$\oint dw\>\langle s(w)\mathcal{O}(z)\cdots\rangle.$ (39)
When the neutrality condition is satisfied, the vertex operator correlation function transform covariantly under conformal transformations. We transform the screening operator to $w=\infty$, and the behavior of the whole vertex operator correlation function is $\sim w^{-2}$. Performing the contour integral at $w=\infty$, we arrive at the desired result 0.
@@
where $P$ is some polynomials of the $\mathcal{W}$ intertwiners. In this work, we omit the $P$ part, and call the exponential part the free-field vertex operators $e^{i(\alpha_{+}\Lambda^{(+)}+\alpha_{-}\Lambda^{(-)})\cdot\phi^{i}}:(z)$. The dual vertex operators are also introduced, with the shift of the $g$ weight as $\Lambda\to-\Lambda-2\alpha_{+}\rho-2\alpha_{-}\rho^{\vee}$. Both the dual vertex operators and vertex operators with $U\Lambda$ will be used to represent charge conjugations in the free-field approach.
The application to disk correlation function calculations have been pioneered in Kawai:2002pz ; Caldeira:2003zz ; Hemming:2004dm . The free-field vertex operator realization of a $\mathcal{W}\big{[}\widehat{g}\big{]}(p,p^{\prime})$ disk bulk $n$-point functions take the form of
$_{\Omega^{\prime}}\langle\langle L_{p}|\prod_{j=1}^{M}\oint s^{\pm}_{i}(z^{(j)})\prod_{\widebar{j}=1}^{\widebar{M}}\oint\widebar{s}^{\pm}_{i}(\widebar{j}^{(j)})\prod_{k=1}^{N}V_{\alpha_{k}}(z_{k})\prod_{\widebar{k}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{k}}(\widebar{z}_{k})|0\rangle$,
@@
$_{\Omega^{\prime}}\langle\langle L_{p}|\prod_{i=1}^{N}V_{\alpha_{i}}(z_{i})\prod_{\widebar{i}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{i}}(\widebar{z}_{i})|0\rangle=\delta\widebar{\delta}\prod_{i<j}z_{ij}^{\alpha_{i}\cdot\alpha_{j}}\times\prod_{\widebar{i}<\widebar{j}}\widebar{z}_{\widebar{i}\widebar{j}}^{\widebar{\alpha}_{\widebar{i}}\cdot\widebar{\alpha}_{\widebar{j}}}\times\prod_{i,\widebar{j}}(1-z_{i}\widebar{z}_{\widebar{j}})^{\alpha_{i}\cdot V_{\Omega^{\prime}}\cdot\widebar{\alpha}_{\widebar{j}}}$. (41)
@@
Techniques in integrals.
We only consider calculations of disk two-point functions, and there are two crucial techniques in the integrations.
@@
  • Consider an integral with $n$ holomorphic screening operators insertions. We deformed the contours into the following integral along segments (with the phase factors absorbed into the normalization)
    $\oint dz^{(1)}\oint dz^{(2)}\cdots\oint dz^{(n)}\longrightarrow\int_{z_{2}}^{z_{1}}dz^{(1)}\int_{z_{2}}^{z^{(1)}}dz^{(2)}\cdots\int_{z_{2}}^{z^{(n-1)}}dz^{(n)}.$ (42)
With the deformation (42), we can apply the Euler-type integral expression and the Taylor expansions of the Lauricella’s hypergeometric function $F_{D}^{(n)}\Big{[}a,b_{1},\cdots,b_{n};c;z_{1},\cdots,z_{n}\Big{]}$, $z_{i}\in\mathbb{C}$ repeatedly to obtain the final result. The Taylor expansion of the $F_{D}^{(n)}$ functions at the origin of $\mathbb{C}^{n}$ is given by
@@
Calculations of disk one-point function don’t require any insertions of $\mathcal{W}$ intertwiners. Consider a minimal model with diagonal and charge conjugation modular invariants. Disk one-point functions with Cardy boundary conditions are

### v4

@@ Abstract
We apply the background charged bosonic free-field approach to the rational principal quantum Drinfeld-Sokolov (QDS) $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ minimal models with boundaries, where $g$ is a bosonic simple Lie algebra. We express their Ishibashi states using the free-field Ishibashi states, and calculate disk two-point correlation functions using the free-field vertex operators with intertwiner insertions. The results of the integral expressions in the free-field approach are obtained by repeatedly applying the Euler-type integral expression and the Taylor expansions of Lauricella’s hypergeometric functions $F_{D}^{(n)}$. We also discuss the background free-field approach to the QDS $\mathcal{W}$ algebras related to a semi-simple Lie algebra $g$, and apply a coset free-field resolution to express diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states.
@@ 1 Introduction
Two-dimensional conformal field theories (CFT2) have infinite-dimensional local conformal symmetries, described by a local conformal algebra $\mathcal{A}\otimes\widebar{\mathcal{A}}$. $\mathcal{A}$ and $\widebar{\mathcal{A}}$ are infinite-dimensional Lie algebras, called the chiral and anti-chiral algebras, respectively Belavin:1984vu ; Moore:1988qv . The infinite-dimensional local conformal symmetries allow us to solve the correlation functions in some theories exactly, and free-field approaches are efficient methods. We claim that a CFT2 admits a free-field approach if the following three conditions are satisfied.
@@
There is a wide range of RCFTs that are known to admit free-field approaches. Two canonical types of theories that admit a free-field approach are
We apply the background-charged free-field approach to rational principal QDS $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ ($g$ being simple Lie algebras), including free-field expressions of minimal Ishibashi states, and detailed example calculations of disk correlation functions using the free-field vertex operators, with insertions of $\mathcal{W}$ intertwiners. The method to calculate the correlation functions is a direct generalization of the celebrated Coulomb gas formalism of the Virasoro minimal models Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi . We show that the integral expressions of free-field vertex operators with $\mathcal{W}$ intertwiners are always solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type of differential equations of the QDS $\mathcal{W}$ minimal models Belavin:1984vu . The crucial techniques in performing the integrals are
@@
Divergent terms can arise in the process. However, due to the fact that the integral expressions are solutions to the BPZ type of differential equations, it is necessary to include the divergent terms. We give sample calculations of disk two-point functions in some of the simplest diagonal and charge-conjugated $\mathcal{W}$ minimal models, including
  • The Virasoro minimal models $\mathfrak{Vir}(5,2)$, $\mathfrak{Vir}(4,3)$, and $\mathfrak{Vir}(5,3)$.
  • The $\mathcal{W}_{3}(5,3)$ and the $\mathcal{W}_{3}(5,4)$ minimal models.
  • The $\mathcal{W}_{4}(7,4)$ model.
  • The $\mathcal{W}\big[\widehat{G}_{2}\big](6,7)$ model.
The calculations of the disk correlation functions with higher points and in other rational $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ models can be obtained from analogous procedures in principle.
In the appendices, we discuss the background-charged free-field approach to the principal QDS $\mathcal{W}$ algebras related to semi-simple Lie algebras. Each simple part of a semi-simple QDS $\mathcal{W}$ algebra decouples, and hence the chiral characters, modular properties, fusion rules, and correlation functions of the semi-simple $\mathcal{W}$ theory are multiplications of those of the related simple $\mathcal{W}$ theories. Finally, we apply a coset Fock space resolution Bouwknegt:1990fb ; Bouwknegt:1990wa to express the diagonal ADE coset $\mathcal{W}$ minimal Ishibashi states Goddard:1984vk ; Goddard:1986ee ; Mathieu:1990dy ; Bouwknegt:1992wg ; Arakawa:2018iyk .
@@ 2.1 Preliminaries
For each element $w\in W(g)$, we can define its length $l(w)$ as the minimum possible number of elements in its simple Weyl reflection decompositions. There is a unique longest element $w_{l}\in W(g)$ and the charge conjugation of $\lambda$ is defined as $U\lambda:=-w_{l}\lambda$. For the $B_{r}$, $C_{r}$, $D_{r}$ even series, $E_{7,8}$, $F_{4}$, and $G_{2}$ the charge conjugation $U=-w_{l}=I$. For the $A_{r}$ series, $U$ reverse the order of all Dynkin labels. For the $D_{r}$ odd series, $U$ reverse the final two Dynkin labels $\lambda_{r-1}\leftrightarrow\lambda_{r}$. For $E_{6}$, $U$ reverse two pairs of Dynkin labels $\lambda_{1}\leftrightarrow\lambda_{5}$ and $\lambda_{2}\leftrightarrow\lambda_{4}$. We define two types of shifted Weyl actions as
@@
Dominant and codominant weight vectors are defined analogously to the finite cases, with the addition of an extra zeroth dimension. The set of level-$k$ dominant and codominant $\widehat{g}(k)$ weight vectors are denoted by $\widehat{P}_{+}^{k}$ and $\widehat{P}_{+}^{\vee k}$ respectively. The principal admissible weight vectors are defined in Frenkel:1992ju . We are only interested in the admissible $\widehat{g}(k)$ weights taking the form of
@@
Define the screening operators of a chiral algebra $\mathcal{A}$ as the free-field vertex operators whose OPEs with free-field chiral generators are regular or a total derivative singular term. From the definition, we conclude that a closed contour integral over an $\mathcal{A}$ screening operator is an $\mathcal{A}$ intertwiner.
The free-field approach to the simple KM $\widehat{g}(k)$ algebras introduced here is the Wakimoto-type, which is realized from $r$ background-charged bosonic fields $\phi^{i}$ and $\big|\Delta_{+}\big|$ bosonic $\beta\gamma$ ghost systems Wakimoto:1986gf ; Bouwknegt:1990aa ; Bouwknegt:1990wa . We present a free-field resolution conjecture for dominant and admissible modules $\widetilde{L}_{\widehat{\Lambda}}$ using the $\phi^{i}\beta^{\alpha}\gamma^{\alpha}$ Fock space modules. Consider a simple KM algebra at positive integer level $k\in\mathbb{Z}_{+}$, where the chiral modules are the irreducible dominant modules $\widetilde{L}_{\widehat{\Lambda}}$, $\widehat{\Lambda}\in\widehat{P}^{k}_{+}$. The screening operators of this theory take the form of Bouwknegt:1990wa
@@
Let $C_{(\Lambda^{(+)},\Lambda^{(-)})}=(F_{(\Lambda^{(+)},\Lambda^{(-)})},d)$ be the BRST image of the $\widehat{g}(k)$ admissible resolution complex $\widetilde{C}_{(\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}$, where $d$ are $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ intertwiners. The BRST image of an admissible (and dominant) $\widehat{g}(k)$ modules is either zero or irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ of the $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ algebras Frenkel:1992ju . Hence, the zeroth cohomology of the complex $(F_{(\Lambda^{(+)},\Lambda^{(-)})},d)$ is either trivial or being isomorphic to the irreducible modules $L_{(\Lambda^{(+)},\Lambda^{(-)})}$ Bouwknegt:1991gf .
@@
$\Big\{L_{(\Lambda^{(+)},\Lambda^{(-)})}\quad|\quad\widehat{\Lambda}^{(+)}\in\widehat{P}^{p-h^{\vee}}_{+},\>\widehat{\Lambda}^{(-)}\in\widehat{P}^{\vee p^{\prime}-h}_{+}\Big\},$ (20)
@@
$(L_{n}-\widebar{L}_{-n})|b\rangle_{\Omega}=\big[W_{n}^{i}-(-1)^{h_{i}}\Omega\widebar{W}^{i}_{-n}\big]|b\rangle_{\Omega}=0.$ (25)
For each $\mathcal{A}$ representation $L_{i}$ (not necessarily irreducible), the solutions to the conditions (25) are the $\Omega$-twisted Ishibashi states $|L_{i}\rangle\rangle_{\Omega}$ Ishibashi:1988kg ; Onogi:1988qk
$|L_{i}\rangle\rangle_{\Omega}=\sum_{N=0}^{\infty}|i,N\rangle\otimes V_{\Omega}U|i,N\rangle,$ (26)
where $|i,N\rangle$ denotes the orthonormal basis of $L_{i}$. $U$ is an anti-unitary isomorphism mapping a representation to its charge conjugation. $V_{\Omega}$ is a unitary isomorphism induced from $\Omega$ such that $V_{\Omega}:\>L_{i}\to L_{\omega^{-1}(i)}$ Recknagel:1997sb ; Recknagel:2013uja .
The boundary states are linear combinations of the Ishibashi states. To obtain the form of the physical boundary states, we need to consider the constraint conditions from two distinct yet equivalent methods of calculating the annulus partition functions Cardy:1989ir . The constraint condition comes from the form of the open-sector Hilbert space between two boundaries of RCFTs
@@
One simple type of solution satisfying the $n_{b_{1}b_{2}}^{i}\in\mathbb{N}$ constraint condition is the Cardy boundary states, where the boundary states have the same labeling as the irreducible modules in the Kac-table, with the celebrated $n=N$ relation Cardy:1989ir . The explicit form of the Cardy boundary states is
$|b_{i}\rangle_{\Omega}=\sum_{j}\frac{S_{ij}}{\sqrt{S_{0j}}}|L_{j}\rangle\rangle_{\Omega}.$ (28)
@@ 2.2 Universal results
$|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{\Omega}=\sum_{\widehat{w}\in\widehat{W}}\theta_{\widehat{w}}|F_{(\widehat{w}\ast\widehat{\Lambda}^{(+)},\widehat{\Lambda}^{(-)})}\rangle\rangle_{\Omega^{\prime}},\quad\Omega=U,I,$ (29)
@@
(33) means that all the $W^{i}$ gluing conditions are satisfied if the Virasoro gluing condition are satisfied, and the anti-holomorphic momentum $\widebar{p}$ provide the correct $\widebar{w}^{i}$ eigenvalue.
@@
We conjecture that $w^{i}$ eigenvalues are the standard symmetric polynomial of order $i$ of the variables $\theta_{j}=(\Lambda+\alpha_{+}\rho+\alpha_{-}\rho^{\vee},\epsilon_{j})$, where $\{\epsilon_{j}\}$ is the set of weight of vector representation of $g$. This conjecture is a direct generalization of the $w^{3}$ eigenvalue shown in Bouwknegt:1992wg . The $w^{i}$ eigenvalues changed under the action $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ as $w^{i}\to(-1)^{h^{i}}w^{i}$. This is the expected the $w^{i}$ eigenvalue shift under the charge conjugation
$U:\quad w^{i}\to(-1)^{h^{i}}w^{i},$ (35)
since the trivial $W^{i}_{0}$ gluing conditions is $\big[W^{i}_{0}-(-1)^{h^{i}}\widebar{W}_{0}\big]|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{I}=0$, with the corresponding solution being and Ishibashi state in the charge conjugated model. Hence, we can understand that the shift $\Lambda\to-\Lambda-\alpha_{+}\rho-\alpha_{-}\rho^{\vee}$ is a form of charge conjugation in the free-field realization.
@@
$\big[W^{i}_{0}-(-1)^{h^{i}}\widebar{W}^{i}_{0}\big]|L_{(\Lambda^{(+)},\Lambda^{(-)})}\rangle\rangle_{I}=0,$ (36)
@@
The method we use to calculate the $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ correlation functions, is a direct generalization of the Coulomb gas formalism of the Virasoro minimal models Dotsenko:1984nm ; Dotsenko:1984ad ; Dotsenko:1985hi , where the sphere correlation functions are calculated from free-field vertex operators with insertions of intertwiners to satisfy the neutrality condition. This method can be understood as the following way: though the free-field vertex operators could provide the correct chiral algebra eigenvalues, the fusion rules of the two theories don’t match. Hence, intertwiners are introduced to map the $\mathcal{W}$ minimal fusion rules to the free-field momentum conservation rules, without bothering the results in the original theory. We prove the validity of this method by showing that the integral expressions obtained are the solutions to the Belavin-Polyakov-Zamolodchikov (BPZ) type differential equations when the neutrality conditions are satisfied Belavin:1984vu .
@@
$\langle\chi(z)\cdots\rangle=D\langle\mathcal{O}(z)\cdots\rangle=0,$ (38)
where $D$ is some differential operator. The chiral algebra $\mathcal{A}$ intertwiners map $\mathcal{A}$ primary vectors to $\mathcal{A}$ primary vectors, hence, any singular vector can be expressed as $|\chi\rangle=Q^{\chi}|v\rangle$, where $Q^{\chi}$ is an intertwiner and $|v\rangle$ is a non-singular primary vector Thorn:1984sn ; Bouwknegt:1992wg . This allow us to replace the singular operator $\chi(z)$ by a commutator $Q^{\chi},\mathcal{O}(z)]$, for some primary operator $\mathcal{O}(z)$, corresponding to $|v\rangle$. Then, by writing the intertwiner $Q^{\chi}$ as (multiple) contour integral over screening operators, we obtain
$\oint dw_{1}\cdots\oint dw_{i}\>\langle s_{1}(w_{1})\cdots s_{i}(w_{i})\mathcal{O}(z)\cdots\rangle.$ (39)
When the neutrality condition is satisfied, the vertex operator correlation function transform covariantly under conformal transformations. We transform the screening operator $s_{1}(w_{1})$ to $w_{1}=\infty$, and the asymptotic behavior of the correlator is $\sim w_{1}^{-2}$. Performing the $w_{1}$ contour integral at the infinity point $\infty$, we arrive at the desired result 0. This procedure showed how singular vectors are obtained from the language of the free-fields and ensured the integral expressions to be the solutions of the BPZ type of differential equations.
@@
where $P$ is some polynomials of the $\mathcal{W}$ intertwiners. In this work, we omit the $P$ part, and call the exponential part the free-field vertex operators $e^{i(\alpha_{+}\Lambda^{(+)}+\alpha_{-}\Lambda^{(-)})\cdot\phi^{i}}:(z)$. We also introduce the so-called dual vertex operators, with the shift of the $g$ weight as $\Lambda\to-\Lambda-2\alpha_{+}\rho-2\alpha_{-}\rho^{\vee}$. Both the dual vertex operators and vertex operators with $U\Lambda$ will be used to represent charge conjugations in the correlation function calculations.
The application of free-field approaches to disk correlation function calculations was pioneered in Kawai:2002pz ; Caldeira:2003zz ; Hemming:2004dm . The free-field vertex operator realization of a $\mathcal{W}\big[\widehat{g}\big](p,p^{\prime})$ disk bulk $n$-point functions take the form of
$_{\Omega}\langle\langle L_{p}|\prod_{j=1}^{M}\oint s^{\pm}_{i}(z^{(j)})\prod_{\widebar{j}=1}^{\widebar{M}}\oint\widebar{s}^{\pm}_{i}(\widebar{j}^{(j)})\prod_{k=1}^{N}V_{\alpha_{k}}(z_{k})\prod_{\widebar{k}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{k}}(\widebar{z}_{k})|0\rangle$,
@@
$_{\Omega}\langle\langle L_{p}|\prod_{i=1}^{N}V_{\alpha_{i}}(z_{i})\prod_{\widebar{i}=1}^{\widebar{N}}\widebar{V}_{\widebar{\alpha}_{i}}(\widebar{z}_{i})|0\rangle=\delta\widebar{\delta}\prod_{i<j}z_{ij}^{\alpha_{i}\cdot\alpha_{j}}\times\prod_{\widebar{i}<\widebar{j}}\widebar{z}_{\widebar{i}\widebar{j}}^{\widebar{\alpha}_{\widebar{i}}\cdot\widebar{\alpha}_{\widebar{j}}}\times\prod_{i,\widebar{j}}(1-z_{i}\widebar{z}_{\widebar{j}})^{\alpha_{i}\cdot V_{\Omega^{\prime}}\cdot\widebar{\alpha}_{\widebar{j}}}$. (41)
@@
Techniques in correlation fucntion calculations.
We only consider calculations of disk two-point functions, and there are two crucial techniques in the calculations.
@@
  • We deformed the contours into the following integral along segments (with the phase factors absorbed into the normalization) between $z_{2}$ and the adjacent branching point. We will determine the ordering of the vertex operators in each detailed calculation, for technical simplicity.
With the contour deformations, we can apply the Euler-type integral expression and the Taylor expansions of the Lauricella’s hypergeometric function $F_{D}^{(n)}\Big[a,b_{1},\cdots,b_{n};c;z_{1},\cdots,z_{n}\Big]$, $z_{i}\in\mathbb{C}$ repeatedly, to obtain the final result. The Taylor expansion of the $F_{D}^{(n)}$ functions at the origin of $\mathbb{C}^{n}$ is given by
@@
Calculations of disk one-point function don’t require any insertions of $\mathcal{W}$ intertwiners. Consider a minimal model with diagonal and charge conjugation modular invariants. Disk one-point functions in $\Omega$ permuted RCFTs, with Cardy boundary conditions are